OpenStudy (anonymous):

find the following limits analytically: f(x)= {x-3, x can't = 1} and {0, x=1} lim x->1 f(x) =

OpenStudy (anonymous):

what does "analytically" mean?

OpenStudy (anonymous):

the limit is \(1-3=-2\) because the limit does not depend on the actual value of the function at \(x=1\) but only on what the function approaches. since you have \(f(x)=x-3\) for all \(x\) except \(x=1\) the limit is -2

OpenStudy (anonymous):

thanks! how do i figure out the second one?

OpenStudy (anonymous):

what second one? i thought this was one question, not two

OpenStudy (anonymous):

its in brackets and is one big question it is x-3, x can't= 1 0, x=1 all in brackets

OpenStudy (anonymous):

yes it is a piecewise function and you are asked for the limit as x goes to 1 answer is -2

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (anonymous):

i have a few more questions to ask if that is ok